# Signal Processing over Time-Varying Graphs: A Systematic Review

### v1

@@ IV Graph Signal Processing for Time-varying Graphs
When dealing with time-varying graph signals, GSP algorithms can be broadly categorized into online and offline methods. Online GSP algorithms, such as adaptive filters and state-space models, update dynamically with time-varying data, offering real-time processing capabilities. Offline methods, on the other hand, analyze graph signals in batches, leveraging the entire dataset to capture spatiotemporal relationships. In this section, we will focus on non-GNN approaches, covering algorithm genres such as adaptive filtering, time-series models, and their extensions for graph-structured data.
IV-A Online GSP Algorithms
@@
Let us begin with the discussion of online GSP algorithms for time-varying graph signals. The online GSP algorithms share a workflow similar to that of the online algorithms found in non-graph contexts. Following the convention seen in most adaptive GSP algorithms, we assume that the data sequence arrives one instance at a time, with each instance will be denoted as . In some cases, the data can be partially observed, in which the partial observation can be modeled by applying a sampling matrix to the receive time instance . The data received can also contain noise, which is often represented as an additive noise term . Typically, there will be a weight term defined by utilizing the graph and serves as how the observation is modified or processed. Without the loss of generality, we denote this weight term as but keep in mind that the type and dimension of this term are algorithm-specific. In many cases, the weight is updated based on calculating a metric between the estimation and the observation . The online GSP algorithm update under noisy and missing data observations can be generalized using the following framework:
@@
where denotes the graph signal update function and denotes the weight update function. An illustration of the workflow of online GSP algorithms is shown in Figure 3.
@@
1) GSP-based (Spectral) Adaptive filters: A natural idea for the online processing of time-varying node signals is to utilize adaptive graph filters that consist of a combination of classical adaptive filters with graph shift operations and can efficiently process time-varying node signals owing to their simplicity of implementation. Similar to the classical adaptive filters, GSP-based adaptive filters update in the direction opposite to the error at each time step based on convex optimization. What makes this combination different from the classical adaptive filters is that most GSP-based adaptive filters utilize a predefined bandlimited filter in the spectral domain derived from GFT, which implies that it captures the topological information of the graph in the spectral domain along with the time-varying signal. Using the framework we defined in (7) and the graph convolution in (4), the weight in GSP-based adaptive filters can be expressed as . By leveraging topological information, GSP-based adaptive filters enable signal reconstruction across both spatial and temporal dimensions. The first GSP-based adaptive filter is the adaptive graph Least Mean Squares (GLMS) algorithm[13]. In the GLMS, the update strategy is the optimization results derived based on a -norm optimization with the assumption of Gaussian noise:
@@
The plugging (8) into (9), update function of GLMS can be expressed as
@@
where is the step-size parameter. There are several variants and extensions of the GLMS. The GLMS was originally proposed on a spatiotemporal graph where the graph signals are time-varying but the topology is static; by introducing small perturbations to the graph Laplacian matrix, the GLMS can be extended to dynamic graphs [60]. A variant of the GLMS is the Graph Normalized LMS (GNLMS), in which a spectral domain normalization is derived from minimizing the spectral error between consecutive estimations and then introduced to the weights of the LMS-based update function to improve the convergence behavior [2]. Another variant is the graph recursive least squares (GRLS), which leverages the RLS framework to provide enhanced convergence behavior and superior prediction performance compared to GLMS [61, 62]. By recursively minimizing a weighted least squares error criterion, GRLS adapts more quickly and more accurately to time-varying changes under noisy signals when compared to GLMS.
@@
Gaussian noise assumption is seen in most noise models, and -norm optimization is the go-to option for Gaussian noise because minimizing the squared error corresponds to the maximum likelihood estimate solution [63]. However, the underlying noise in a variety of realistic applications, including meteorological recordings [64] and powerline communication [65], is verified to possess impulsive behaviors that could be represented by heavy-tailed, non-Gaussian distributions, such as generalized Gaussian, Student’s t, and -stable distributions [63]. The least-squares-based methods lose their validity in impulsive noise situations owing to the presence of large or infinite variance caused by the impulsiveness in the noise [66]. Several approaches can be taken to overcome this drawback. Instead of using second power, a least mean power is used to define a loss based on the assumption that the noise is symmetric -stable distribution (SS), giving us the Graph Least Mean algorithm [67]. The Normalized GLMP was further proposed to include a time-varying spectral normalization matrix to speed up the computation of GLMP [68]. Moving further with the idea of -norm optimization, the adaptive Graph-Sign algorithm (Graph-Sign) was derived as a graph extension of the classical adaptive sign error or the least mean absolute deviation (LMAD) algorithm for multivariate signals [69, 70]. The Graph-Sign algorithm is derived from the minimum dispersion criterion and then reduced to -norm optimization, which removes the need for prior knowledge from any noise assumption [71]. This allows the Graph-Sign algorithm to avoid the instability seen in the least-squares-based algorithms when estimating graph signals under impulsive noise.
@@
with and being the regularization term. Taking GLMP as a specific example, the GLMP update function following the framework shown in (7) can be obtained by plugging (11) into (9):
where is the Hadamard product and the superscript p-1 denotes element wise power.
@@
An alternative approach than simply modifying the loss function by switching -norm to -norm or -nor to enhance robustness is to employ in which a robust adaptive estimation of graph signals is developed by replacing the conventional of using -norm to define optimization problem with the Welsch loss [72]. The Welsch loss differs from conventional loss functions by incorporating robustness as a continuous parameter, allowing the Welsch loss to adapt dynamically during algorithm deployment [73]. Instead of directly changing the cost function, a less extreme method to the problem is to add a regularization term to the LMS cost function: the Graph Alternating Minimization algorithm improves the robustness of GLMS by separately modeling the impact of Gaussian noise and outliers in the cost function [74]. The above GSP-based adaptive filters can be further generalized to have the cost function
@@
2) GSP-based (Spatial) Adaptive filters: Notice that the adaptive GSP algorithms are not limited to the spectral domain but could also be conducted in the spatial domain by using the Chebyshev polynomial approximation shown in (6). For example, the bandlimited filters in the GLMS can be approximated using a series of Chebyshev polynomials, leading to the spatial domain graph diffusion [75]. A GSP-based adaptive filter that uses a combination of graph diffusion and sign update strategy was recently proposed in [76]. Recently, the Graph Signal Adaptive Message Passing has been proposed, which is an adaptive filter defined using the flexible localized node message passing scheme instead of defining all computations globally on the graph level [77]. In general, we can use the Chebyshev polynomial approximation shown in (6) to define the weight in the cost function such as the ones seen in (8) and (11):
@@
where are the coefficients for after merging the terms within . It is worth mentioning that the spatial methods resemble the structure of many GNNs, as the majority of GNNs are derived based on the (spatial) GCN [8, 10]. In short, one layer of GCN can be obtained by some algebraic manipulations of (6) and then feeding its output into an activation function. We will be conducting a more thorough review of GNN-based methods for time-varying graph signals in Section V.
@@
3) State Space Models for TV-GSP: Aside from GSP-based adaptive filters, another direction to approach the problem is to use state space models, and the typical example is the Kalman Filter [78]. The Kalman Filter is a recursive algorithm used to estimate the state of a linear dynamic system under the Gaussian noise assumption. In GSP, we can formulate system equations suitable for time-varying graph signals by slightly modifying the framework in (7):
@@
The operations within and can now be conducted entirely in the spectral domain using spectral graph convolution in (4). With the multi-variate signals represented as graph signals, the rest of the procedure, such as the update step and the prediction step, follows the classical Kalman filter workflow, with adjustments for the graph representation. Once the estimation and update steps are completed, the results are transformed back into the spatial domain using the Inverse Graph Fourier Transform (IGFT) in (3).
@@
The Kalman Filter over signals on sensor networks can be established for which the optimal estimation method at each sensor is done when the sensor communications are formed by an undirected graph [80]. The Extended Kalman filter (EKF) extends the applicability of the standard Kalman filter from linear systems to non-linear systems by locally approximating them as linear [81]. The graph version of the Extended Kalman filter is a graph process obeying a nonlinear state-space model that can be found by forming a GSP-based Bayesian filter whose goal is to minimize the computed covariance matrix as the Kalman gain of the Extended Kalman filter in the graph spectral domain [79]. The Unscented Kalman Filter (UKF) addresses the non-linearity of systems by representing the mean and covariance of the state using a set of deterministically chosen sample points to avoid the need for linearization, which is different from the standard Kalman filter [82]. This approach can also be applied to graph signals, which leads to the Unscented Kalman filter of graph signals by again using the decomposition of graph Laplacian matrix to form the GFT and then design the Kalman gain matrix from the eigenvectors [83]. However, the Kalman filter is designed for Gaussian noise because of minimizes the mean square error of the estimated states, which is not suitable for non-Gaussian noise.
@@
Remarks Online algorithms are most preferred when the outputs are required to be made on the fly as new inputs are received. Their ability to produce real-time outputs based on very limited data observations in the temporal domain, without requiring the entire input upfront, makes them highly effective in TV-GSP tasks due to the fact that most time-varying graph signals already have a large spatial dimension. Online GSP algorithms are not limited to regression tasks; their outputs can also serve as inputs for classification or clustering models [31, 84]. This capability broadens the application scope of online GSP algorithms. On one hand, online algorithms often have lower computational complexity to cope with real-time output capability, which is a critical advantage of deploying online algorithms on time-varying graph signals. On the other hand, this same online requirement often restricts the length of the prediction horizon of online algorithms, limiting their ability to generate long-term temporal predictions. In contrast, offline algorithms process the entire dataset at once, enabling long-term predictions on graph signals. In the next subsection, we will be discussing offline GSP-based algorithms for time-varying graph signals.
IV-B Offline GSP Algorithms

### v2

@@ 3 Signal Processing for Time-varying Graphs
When dealing with TVG signals, GSP algorithms can be broadly categorized into online and offline methods. Online TVGSP algorithms, such as adaptive filters and state-space models, update dynamically with time-varying data, offering real-time processing capabilities. Offline methods, on the other hand, analyze graph signals in batches, leveraging the entire dataset to capture spatiotemporal relationships. In this section, we will focus on non-GNN approaches, covering algorithm genres such as adaptive filtering, time-series models, and their extensions for graph-structured data.
3.1 Online TVGSP Algorithms
@@
Let us begin with the discussion of online TVGSP algorithms for TVG signals. The online TVGSP algorithms share a workflow similar to that of the online algorithms found in non-graph contexts. Following the convention seen in most adaptive GSP algorithms, we assume that the data sequence arrives one instance at a time, with each instance will be denoted as . In some cases, the data can be partially observed, in which the partial observation can be modeled by applying a sampling matrix to the received time instance . The data received can also contain noise, which is often represented as an additive noise term . Typically, there will be a weight term defined by utilizing the graph , which serves as how the observation is modified or processed. Without the loss of generality, we denote this weight term as , but keep in mind that the type and dimension of this term are algorithm-specific. In many cases, the weight is updated based on calculating a metric between the estimation and the observation . The online TVGSP algorithm update under noisy and missing data observations can be generalized using the following framework:
@@
where denotes the graph signal update function and denotes the weight update function. An illustration of the workflow of online TVGSP algorithms is shown in Fig. 4.
3.1.1 Spectral Adaptive Filters
@@
A natural idea for the online processing of time-varying node signals is to utilize adaptive graph filters that consist of a combination of classical adaptive filters with graph shift operations and can efficiently process time-varying node signals owing to their simplicity of implementation. Similar to the classical adaptive filters, GSP-based adaptive filters update in the direction opposite to the error at each time step based on convex optimization. What makes this combination different from the classical adaptive filters is that most GSP-based adaptive filters utilize a predefined bandlimited filter in the spectral domain derived from GFT, which implies that it captures the topological information of the graph in the spectral domain along with the time-varying signal. Using the framework we defined in (9) and the graph convolution in (4), the weight in GSP-based adaptive filters on STGs can be expressed as . By leveraging topological information, GSP-based adaptive filters enable signal reconstruction across both spatial and temporal dimensions. The first GSP-based adaptive filter is the adaptive graph Least Mean Squares (GLMS) algorithm[31]. In the GLMS, the update strategy is the optimization results derived based on an -norm minimization with the assumption of Gaussian noise:
@@
The plugging (10) into (11), update function of GLMS is
@@
where is the step-size parameter. There are several variants and extensions of the GLMS. By introducing small perturbations to the Laplacian matrix, the GLMS can be extended to dynamic graphs [36]. A variant of GLMS is Graph Normalized LMS (GNLMS), in which a spectral domain normalization is derived from minimizing the spectral error between consecutive estimations and then introduced to the weights of the LMS-based update function to improve the convergence behavior [170]. Graph Recursive Least Squares extends the RLS framework to enhance convergence and prediction under noisy, time-varying signals by recursively minimizing a weighted error criterion, outperforming GLMS [35, 123].
@@
The Gaussian noise assumption is seen in most noise models, and -norm minimization is the go-to option for Gaussian noise because minimizing the squared error corresponds to the maximum likelihood estimate solution [23]. However, the underlying noise in a variety of realistic applications, including meteorological recordings [15] and powerline communication [84], is verified to possess impulsive behaviors that could be represented by heavy-tailed, non-Gaussian distributions, such as generalized Gaussian, Student’s t, and -stable distributions [23]. The least-squares-based methods lose their validity in impulsive noise situations owing to the presence of large or infinite variance caused by the impulsiveness in the noise [92]. Several approaches can be taken to overcome this drawback. Instead of using the second power, the least mean power is used to define a loss based on the assumption that the noise is a symmetric -stable distribution (SS), giving us the Graph Least Mean algorithm [129]:
@@
with and being the regularization term. The GLMP update function following the framework shown in (9) can be obtained by plugging (13) into (11):
where is the Hadamard product and the superscript p-1 denotes element wise power. The Normalized GLMP was further proposed to include a time-varying spectral normalization matrix to speed up the computation of GLMP [196]. Moving further with the idea of -norm minimization, the adaptive Graph-Sign algorithm (Graph-Sign) was derived as a graph extension of the classical adaptive sign error or the least mean absolute deviation (LMAD) algorithm for multivariate signals [197, 164]. The Graph-Sign algorithm is derived from the minimum dispersion criterion and then reduced to -norm minimization, which removes the need for prior knowledge from any noise assumption [188]. The resulting update algorithm is (9) can be obtained by plugging (13) into (11):
Because the Graph-Sign is obtained from an -norm problem, the resulting update is robust. To view it from another angle, the Sign function is a clipping function, which is not affected by the outliers and extreme values in the error , making the Graph-Sign robust. This allows the Graph-Sign algorithm to avoid the instability seen in the least-squares-based algorithms when estimating graph signals under impulsive noise.
@@
An alternative approach to simply modifying the loss function by switching -norm to -norm or -norm to enhance robustness is to employ a robust adaptive estimation of graph signals developed by replacing the convention of using -norm to define the optimization problem with the Welsch loss [189]. The Welsch loss differs from conventional loss functions by incorporating robustness as a continuous parameter, allowing the Welsch loss to adapt dynamically during algorithm deployment [7]. Instead of directly changing the cost function, a less extreme method to the problem is to add a regularization term to the LMS cost function. The Graph Alternating Minimization algorithm improves the robustness of GLMS by separately modeling the impact of Gaussian noise and outliers in the cost function [101]. The graph median filter has also shown good performance under impulse noise [176, 177].
Extending adaptive filters to DTDGs and CTDGs: The adaptive GSP filters reviewed here were initially proposed for STGs; however, their formulation is readily generalizable to DTDGs by introducing time dependency into the graph operators and filter coefficients. At present, however, to the best of our knowledge, this generalization has not been systematically formalized in the literature, and existing work rarely addresses adaptive filtering explicitly in dynamic graph settings. Our discussion here highlights an important and largely unexplored research direction, where the adaptive TVGSP framework could be extended to support time-evolving graphs.
Given a DTDG with graph at time , to define the GSP-based adaptive filters, we can replace in (11) with
where is obtained from the GFT of , which is from the eigenvector decomposition of the instantaneous graph Laplacian matrix .
Alternatively, instead of defining the spectral operator directly from the instantaneous Laplacian at each time, we can view the DTDG as producing incremental changes with respect to the previous graph. Let
denote the adaptive GSP filter at time . When the graph evolves to , we write
where captures the effect of the topology change, for example via a change of the GFT basis,
and the change of the operator
To provide a demonstration using GLMS in (12), the DTDG version can be expressed as
Similarly, the DTDG version of GLMP is
and the DTDG version of Graph-Sign is
For the CTDG counterpart, it is more natural to model changes as event-driven rather than discrete temporal updates. In this case, instead of indexing the adaptive GSP filter by discrete time instants, we associate each update with an event that modifies the graph topology. Let
denote the adaptive spectral filter immediately before the event occurs. When the graph evolves due to event , the updated filter can be written as
where captures the spectral effect of the event-driven topology change. This perturbation may arise from a change in the GFT basis,
and a corresponding update of the spectral operator,
This event-based formulation highlights how CTDG evolution can be incorporated into adaptive GSP filtering by explicitly linking filter updates to discrete structural events rather than uniform temporal sampling.
Compared with the STG case, extending adaptive spectral GSP methods to DTDGs and CTDGs requires recomputing the GFT whenever the topology of changes, resulting in a computational burden on the order of over each evolving graph. This becomes challenging for large-scale or rapidly evolving systems, and repeated eigendecompositions may also introduce numerical instability for large . Currently, extensions of TVGSP to DTDGs and CTDGs remain limited due to this scalability issue, with existing approaches such as Sardellitti et al. [160] where topology changes through small perturbations that allow only 1 edge update per topology change. On the other hand, these challenges highlight a compelling opportunity for developing more stable, efficient, and theoretically grounded graph operators tailored to dynamic graph settings. Designing such scalable GFT surrogates and adaptive spectral frameworks for TVGs remains an exciting and largely open research direction.
3.1.2 Spatial Adaptive Filters
@@
Notice that the adaptive GSP algorithms are not limited to the spectral domain but could also be conducted in the spatial domain by using the Chebyshev polynomial approximation shown in (7). For example, the bandlimited filters in the GLMS can be approximated using a series of Chebyshev polynomials, leading to the spatial domain graph diffusion [124]. A GSP-based adaptive filter that uses a combination of graph diffusion and sign update strategy was recently proposed in [198]. Recently, the Graph Signal Adaptive Message Passing has been proposed using the flexible localized node message passing scheme instead of defining all computations globally on the graph level [200]. In general, we can use the Chebyshev polynomial approximation shown in (7) to define the weight in the cost function, such as the ones seen in (10) and (13):
@@
where are the coefficients for after merging the terms within . It is worth mentioning that the spatial methods resemble the structure of (spatial) GCN-based GNNs: a GCN layer is obtained by algebraic manipulations of (7), then feeding its output into an activation function [86, 34]. We will conduct a more thorough review of GNN-based methods for TVG signals in Section 4.
We can again define the DTDG versions of the GSP-based spatial adaptive filters using logic similar to how the DTDG GSP-based (spectral) adaptive filters are defined. In the case of the spatial algorithms, for a DTDG , all the operations are directly defined by the graph Laplacian matrix , meaning that we replace in (28) with
Alternatively, instead of constructing the spatial operator directly from the instantaneous Laplacian , one may interpret the evolution of the DTDG as producing incremental changes with respect to the operator at the previous time step. Let
denote the spatial GSP adaptive filter at time . When the graph evolves to , we write
where accounts for the effect of the topology variation. For instance, this can be modeled via a temporal-modification of the Laplacian,
and the corresponding variation of the filter coefficients
Similar to the spectral case, the CTDG version of spatial adaptive filters has not yet been fully realized, largely due to the difficulty of event-driven graph updates within a stable GSP representation framework. Recall that spatial filtering can be related to spectral filtering through the Chebyshev polynomial approximation in (7), which effectively serves as a bridge between spectral operators and spatial-domain implementations. Building on this connection, one could follow a similar line of reasoning as in the DTDG case: an event-driven formulation could then be derived from (24) to a corresponding spatial adaptive operator in (27).
3.1.3 State Space Models for TVGSP
@@
Aside from GSP-based adaptive filters, another direction to approach the problem is to use state space models, and the typical example is the Kalman Filter [83]. The Kalman Filter is a recursive algorithm used to estimate the state of a linear dynamic system under the Gaussian noise assumption. In GSP, we can formulate system equations suitable for TVG signals by slightly modifying the framework in (9):
@@
The operations within and can now be conducted entirely in the spectral domain using spectral graph convolution in (4). With the multivariate signals represented as graph signals, the rest of the procedure, such as the update step and the prediction step, follows the classical Kalman filter workflow, with adjustments for the graph representation. Once the estimation and update steps are completed, the results are transformed back into the spatial domain using the IGFT in (3).
@@
The Kalman Filter over signals on sensor networks can be established, for which the optimal estimation method at each sensor is done when the sensor communications are formed by an undirected graph [166]. The simplicial complex version of Kalman filters on graphs is also realized through the Hodge Laplacian representation [121]. The Extended Kalman filter (EKF) extends the applicability of the standard Kalman filter from linear systems to non-linear systems by locally approximating them as linear [143]. The graph version of the Extended Kalman filter is a graph process obeying a nonlinear state-space model that can be found by forming a GSP-based Bayesian filter whose goal is to minimize the computed covariance matrix as the Kalman gain of the Extended Kalman filter in the graph spectral domain [155]. Recent work further extends EKF to incorporate graph topology information via the graph Laplacian, enabling more robust localization in networked dynamic systems. For instance, in the context of connected and automated vehicles, the graph Laplacian Extended Kalman Filter leverages the structural relations between vehicles to enhance estimation accuracy and stability under nonlinear dynamics [137]. The Unscented Kalman Filter (UKF) addresses the non-linearity of systems by representing the mean and covariance of the state using a set of deterministically chosen sample points to avoid the need for linearization, which is different from the standard Kalman filter [184]. This approach can also be applied to graph signals, which leads to the Unscented Kalman filter of graph signals by again using the decomposition of the graph Laplacian matrix to form the GFT and then designing the Kalman gain matrix from the eigenvectors [100]. The Kalman filter is designed for Gaussian noise because of minimizes the mean square error of the estimated states, which is not suitable for non-Gaussian noise, making this research challenge to be addressed in the future.
Beyond the Kalman filtering framework, particle filters provide a more general approach capable of handling non-Gaussian and highly nonlinear systems. When extended to graph settings, graph particle filters utilize the underlying topology to improve estimation. In particular, the graph Laplacian distributed particle filter promotes smoothness of state estimates across the network by regularizing neighboring particle estimates based on the graph structure [141]. To further address communication efficiency in distributed implementations, graph-based compression schemes have been proposed. These methods transform local particle distributions into the graph frequency domain to reduce communication overhead while preserving estimation accuracy [208]. Together, these advancements illustrate how graph structure can enhance both Kalman and particle filtering methods, offering promising directions for state estimation in complex networked systems.
Similar to GSP-based adaptive filters, the DTDG and CTDG versions of state-space models for TVGSP are limited. Here, we outline a general approach to realizing state-space models for DTDGs. Topology-based parameters such as can be replaced with their time-varying counterparts to account for the topology changes induced to incorporate the topology change caused by . However, we will not proceed with detailed derivations, proofs, or justifications at this stage, as merging topological changes with the inherently dynamic parameter settings of state-space models poses significant challenges and is left for future research. For CTDGs, the additional complexity of integrating events with GSP and state-space models presents an even greater challenge, requiring dedicated exploration for future research.
@@
Remarks Online algorithms are most preferred when the outputs are required to be made on the fly as new inputs are received. Their ability to produce real-time outputs based on very limited data observations in the temporal domain, without requiring the entire input upfront, makes them highly effective in TVGSP tasks due to the fact that most TVG signals already have a large spatial dimension. Online TVGSP algorithms are not limited to regression tasks; their outputs can also serve as inputs for classification or clustering models [154, 39]. This capability broadens the application scope of online TVGSP algorithms. On one hand, online algorithms often have lower computational complexity to cope with real-time output capability, which is a critical advantage of deploying online algorithms on TVG signals. On the other hand, this same online requirement often restricts the length of the prediction horizon of online algorithms, limiting their ability to generate long-term temporal predictions. In contrast, offline algorithms process the entire dataset at once, enabling long-term predictions on graph signals. In the next subsection, we will be discussing offline GSP-based algorithms for TVG signals.
3.2 Offline GSP Algorithms
3.2.1 General Offline GSP